Demand Qd = 17 - 0.1Pd, Supply = -12 + 0.5Ps.Now suppose that the government establishes a $10 subsidy per each box of tobacco
produced to support domestic producers in the industry against world competitors.
Show how this policy affects the market equilibrium. Calculate the new “consumer
price” and “produce price.”
1
Expert's answer
2020-03-16T12:03:59-0400

After the introduction of a subsidy of $ 10, the equilibrium price will decrease and the volume of sales and sales revenue will increase.

After the introduction of a subsidy of $10 per unit of output, the proposal will change: Qş (P)S→ QS "(P+10);"

"17-0.1P=-12+0.5(P+10);"

"17-0.1P=-12+0.5P+5;"

"0.6P=24"

P"=" $ "40;"

Consumer price"=" $ "40;"

Produce price"=" "50+10=" $ "50;"